OpenStudy (anonymous):

The product of two consecutive positive even integers is 12,320. What are the two integers?

OpenStudy (foolaroundmath):

Let the even numbers be 2m and 2m+2 2m*(2m+2) = 12320 => 4n(m+1) = 12320 => m(m+1) = 3080 now 3080 = 55*56 Thus, m = 55, m+1 = 56 Thus the two even numbers are 2*55 and 2*55 + 2 i.e. 110 and 112

OpenStudy (anonymous):

See, first thing to be noted is Even Integer has a difference of 2, i.e, 2, 4, 6 like that.. Let the first even integer be x and the next consecutive even integer will be: x + 2. So, their product is given as 12,320.. So, x(x + 2) = 12320 \[x^2 + 2x -12320 = 0\] \[x^2 + 112x - 110x - 12320 = 0\] \[x(x+112) - 110(x + 112) = 0\] \[(x-110)(x + 112) = 0\] So, x = 110 or x = -112.. As, they are positive even integers, so, we neglect x = -112.. x = 110 and (x + 2) = 110 + 2 = 112 So, the two numbers are : 110 and 112...